# Epistemological Bias As a Means for the Automated Detection of Injustices in Text

### v1

@@
Abstract
Injustice occurs when someone experiences unfair treatment or their rights are violated and is often due to the presence of implicit biases and prejudice such as stereotypes. The automated identification of injustice in text has received little attention, due in part to the fact that underlying implicit biases or stereotypes are rarely explicitly stated and that instances often occur unconsciously due to the pervasive nature of prejudice in society. Here, we describe a novel framework that combines the use of a fine-tuned BERT-based bias detection model, two stereotype detection models, and a lexicon-based approach to show that epistemological biases (i.e., words, which presupposes, entails, asserts, hedges, or boosts text to erode or assert a person’s capacity as a knower) can assist with the automatic detection of injustice in text. The news media has many instances of injustice (i.e. discriminatory narratives), thus it is our use case here. We conduct and discuss an empirical qualitative research study which shows how the framework can be applied to detect injustices, even at higher volumes of data.
1 Introduction
@@
The most basic duty of the media is knowledge sharing. Yet, the tool necessary for wide-spread knowledge sharing is influence. With this influence, the media is able to shape how one will understand the intricacy of a story-line in a news story with little effort. This often results in the use of epistemological biases which involves propositions that are presupposed, entailed, asserted, hedged, or boosted in text (Recasens et al., 2013) to erode or assert a person’s capacity as a knower, leading to framing issues and injustice within the text. Particularly, some of these word choices in text can lead to injustices occurring. These injustices can affirm or perpetuate stereotypes concerning the subject. Though it has always been a harsh reality with various ringing consequences, in recent years we have publicly witnessed how the affirmation of stereotypes can lead to physical violence, prejudice, and negative self-image (Harrison and Esqueda, 1999; Gover et al., 2020; Kuykendall, 1989). These experiences are harmful and dangerous, explicitly for the victims but for all members of our society.
@@
Therefore, we seek to detect instances of testimonial injustice, what we define as character injustice, and framing injustice. Testimonial injustice occurs when word choices cast modified believability on a statement due to prejudices (e.g. stereotypes) about the sayer (Fricker, 2007). Character assassination is “the deliberate destruction of an individual’s reputation or credibility" (Icks et al., 2019). This often leads to character injustice, which is an unjustified attack on a person’s character that results in an unfair criticism or inaccurate representation of them (e.g. exaggeration or defamation). According to (Entman, 2007) framing bias happens when the use of subjective, one-sided words that reveals the stance of an author occurs. Which means that an individual’s choice from a set of options is influenced more by how the information is worded, rather than by the information itself. We recognise that news content can be positively or negatively framed to influence the narratives. In this paper, we aim to show how word choices which negatively frame a statement - opposed to speaking about it factually or neutrally, when it affects particular subjects or individuals, can lead to framing injustice (e.g. vilification).
@@
With this work, we seek to make room for subjects of text, even in creative writing, to not have their credibility shot or character assumed due to well-known stereotypes which are harmful and unfounded. It can be noted here that toxic or hate speech is often blatant and highly recognisable, here we focus on implicit biases which are less obvious but are just as harmful. We pursue this, acknowledging it is achievable to speak about any happening in a factual way that does not seek to cause harm on the subject of the text.
@@
Our proposed framework will help with the detection of testimonial, character, and framing injustices. The framework includes a fine-tuned BERT model based on work from (Pryzant et al., 2020b) to automatically tag words associated with epistemological bias from an input text, use the (Kwon and Gopalan, 2021) CO-STAR model and (Sap et al., 2020) Social Bias Frames (SBF) to generate potential associated stereotypes and the concepts of those stereotypes to the input text, and show when the tagged words (associated with some epistemological bias) or less credibility of a person are correlated with a stereotype which causes injustice. Though we could use examples from various fields (e.g. politics, marketing, medicine, etc…), we will use news media as a use-case throughout this work. Thus, we present the following contributions: (1) We develop a novel framework that uses the results of 3 models to automatically detect character, testimonial, and framing injustices in News Media, (2) We produce a fine-tuned tagger model to automatically detect epistemological bias, (3) We develop an Interactive User-Interface for journalists and editors to submit text to and receive output and explanations surrounding the tagged word, and (4) We produce empirical evidence showing how epistemological bias can translate to injustices.
@@
An outcome of this work is to give journalists and editors a tool which will help them easily and quickly notice and avoid testimonial, character, and framing injustices in their work. This will be accomplished by showing users which words they use that produce epistemological bias, the potential stereotype associated with the tagged words and text, offering the user explainability with the help of the stereotype concepts as defined by (Kwon and Gopalan, 2021), and resources to reference literature on the particular epistemological bias type(s) identified in their input text.
2 Background
@@
Many works have established it is difficult for the common person to identify a biased word in a sentence and establish the need for computational agents to take on this charge. Section 2.1, 4.1, and 6.1 and Table 8 of (Pryzant et al., 2020b) shows humans have low ability to detect bias and show humans perform worse than their detection model. Recasens et al. (2013) shows in Table 4 of their paper - that the accuracy of Humans annotators on AMT (amazon mechanical turk) was not more than 37.39% for a single detected biased word. The difficulty arises due to us holding our own biases as facts and lack of education on sentence construction. It has also been observed that when human, expert annotators follow guides that help them detect bias, can more easily find instances of such biases. For example, the editors of the Wiki Neutrality Corpus had to follow the Netural Point of View (NPOV) Guidelines  [^1] https://en.wikipedia.org/wiki/Wikipedia:Neutral_point_of_view.
@@
Great efforts have been put towards identifying potential words in text materials that could encourage epistemological bias (Recasens et al., 2013; Hube and Fetahu, 2019; Pryzant et al., 2020b). Following on from the work of Pryzant et al. (2020b), we fine-tuned their tagger model to automatically tag words associated with epistemological bias from an input text. Identifying words which cause epistemological bias is a step towards awareness of social harms. This begs the question, what do we do with our new found knowledge and awareness? What kinds of implications does our use of these words impact society? What communities are affected by these word choices? These are the questions we explore in this work.
@@
Authors Kwon and Gopalan (2021) have trained a model to detect widely-known stereotypes and the concept of those stereotypes in text materials. We leverage the results of their CO-STAR model and the SBF model (Sap et al., 2020) to offer some explainability of the word choices by the model. Associating a particular text with a stereotype and the concept of that stereotype is a critical step towards awareness of social harms that might cause character or testimonial injustice to a particular individual or group. Lack of identifying the words in a sentence which imply and promote these stereotypes leaves us with the undirected burden and question of: how can we address these harms? This will be further discussed in the methods section of the paper.
@@
Raza et al. (2022) developed a pipeline which takes in news articles, detects and masks words that are biased, and suggests words with more neutral text. Whilst the pipeline and library designed by the authors are very good and useful, they however do not consider the linguistic and epistemological bias features as discussed in Recasens et al. (2013) and used by Pryzant et al. (2020b). Unlike our work discussed in this paper, their pipeline produces no way of distinguishing and highlighting epistemological bias types of any identified potential biased word in the sentence. Also, their framework does not attempt to relate potential stereotypes and stereotype concepts that might be associated with causing injustice to a person or group.
3 Methodology
Hamborg et al. (2019) concluded that various forms of media bias is already analysed in the Social Sciences field and can be implemented in an automated fashion by primarily using Natural Language Processing (NLP) techniques. We leverage research methods using NLP and deep learning whilst also using analysis concepts by researchers from social sciences. Thus, we propose a novel, technical framework (Figure 1) which includes using NLP models for detecting potential epistemological biased words and potential stereotypes along with their concepts, which we semantically link to given sentence(s). From the literature, we found very few models which consider linguistic, epistemological features, or common sense reasoning to detect implicit biases or reasoning about potential stereotypes. This in turn influenced our decision to use three baseline models with promising results (i.e. Tagger Model, Co-Star, and SBF).
Figure 1: Testimonial Injustice Technical Framework
@@ 3 Methodology
3.1 Tagger model
@@
The tagger model (see Figure 7 in Appendix A) takes as input sentence(s) (e.g. a headline text, a collection of headlines, etc…), and predicts the probability of each word in the sentence(s) been biased (see Equation 1 and Figure 1). Then it returns the word with the highest probability as the tagged biased word.
$P_{i}=\sigma(b_{i}W^{b}+e_{i}W^{e}+b)$ (1)
Where $b_{i}\in R^{b}$ is a word’s $w_{i}^{s}$ semantic meaning and $e_{i}$ are the experts features as proposed by Recasens et al. (2013) and based on Equation 2.
$e_{i}=R_{e}LU(f_{i}W^{in})$ (2)
The detection model itself is a BERT-based neural sequence tagging model that has been fine-tuned to include the expert linguistic and epistemological bias features from Recasens et al. (2013). An example of the expert feature as listed in Table 3 of Recasens et al. (2013) is the part of speech (POS) tag of each word in the sentence. Another example of an expert feature from vetted experts in linguistics is “assertive verbs”. The model leverages the semantic meaning of each word in the given sentence via the BERT (Kenton and Toutanova, 2019) contextualised word vector.
When there are words in a sentence which do not help with understanding, but are excess information these are known to cloud the judgment of the reader while simultaneously giving the reader reason to be more confident in the conclusions they make (Lynch, 2023; Spira, 2011). We hypothesised here that - the too much information (TMI) can be considered from an English linguistic standpoint to be a sentence instance which has more than 2 descriptive words (i.e. adjectives and adverbs) in it that do not add to its understanding but seeks to cloud the judgement of the readers. We manually analysed a few biased and non-biased headlines and saw an excessive use of descriptors (i.e. adjectives, adverbs) in headlines that were biased. We extend the detection model further by including a feature on whether or not the sentence contained TMI. We created methods in python which makes use of the core NLP dependency tree and a tree traversal algorithm to go through sentences starting from the root node, then count the number of adjectives and adverbs in the sentence to determine if there is "TMI" or "no TMI" based on the hypothesis. TMI is not a known indicator of epistemological bias, but we acknowledge it might contribute to causing doubt which leads to injustice. This is why we introduce this new feature to see if it will be a contributing feature for the tagger model to detect epistemological biases. We conducted an ablation study to see the effects of adding this feature, whilst training the tagger model in Section 4.1.
3.2 CO-STAR Model and Social Bias Frames
@@
Since testimonial injustice occurs because of widely known stereotypes, the outputs of these models will help us inform our users of any potential stereotypes that are exacerbated by their inputted text. The potential stereotype can also be used to determine if a person’s character has been unjustly targeted. This is because stereotypes are not based on the actual truth of the particular person they are attached to and stigmatises the individual. Their presence amidst character assassination is evidence of character injustice. Stereotype concepts will offer explainability as to why a certain word was tagged as being epistemologically biased.
We submit news article headlines to the CO-STAR and SBF models and receive an output of 6 potential stereotypes and 3 stereotype concepts the sentence is related to. Semantic similarity describes how closely related two items (e.g. 2 words or 2 sentences) are in terms of meaning. We use semantic similarity to determine how closely related each of the generated stereotype sentences in the list are to the original headline sentence. After obtaining the vector embedding for the sentences, we used a distance metric on the encoded vectors (cosine similarity metric from sentence transformers), to get the distance scores, then we ranked the outputs based on their semantic similarity to the headlines. The closest stereotype and stereotype concept to the sentence is output to the user as the potential stereotype and concept which casts doubt on the subject of the inputted text.
3.3 Lexicon Lookup
Once the tagger model has returned the top tagged word, we proceed to automatically look up and semantically search the tagged word in the epistemological lexicons from the social sciences - to discover the epistemological bias types it is associated with. These lexicons are from the collection of datasets we discuss in the dataset section 3.5. When a tagged word is not found in the lexicons, we lemmatize said word (considering its context) to find its base word and find the stem word (removing the prefixes and suffixes). We then search for the lemmatized and stemmed words in the lexicons.
3.4 Interactive Interface for Learning about Bias
@@
During the analysis, we first leverage the nltk python library to remove any stop words (i.e. a, the, etc..). We also check the sentence length. We make an assumption here, that if the sentence is less than 3 words, then there is not enough context to analyse the text for potential stereotypes. However, if we have enough context, the content is sent to the epistemological tagger model to find the word with the highest probability of bias (see Figure 6). The conditions from the framework (Figure 1) are then implemented.
In Figure 6 in Appendix A and 2 show screenshots of the the UI after a user has submitted a text for analyzing. The word suspected of causing injustice is highlighted with the certainty score, an explanation of why the word was tagged. When the user clicks Show Details they will see more details as displayed in Figure 2.
Figure 2: UI showing the tagged words, their associated probability for bias, the potential stereotype and stereotype concept and some explainable data created from the lexicon search.
We display the types of associated epistemological bias detected to the user and provide links to resources which explains more on the specific types of biases as shown in Figure 2. See examples of these outputs in table 5 and table 4 of the Appendix.
3.5 Dataset source used

### v2

# Epistemic Bias as a Means for the Automated Detection of Injustices in Text

@@
Abstract.
Injustices in text are often subtle since implicit biases or stereotypes frequently operate unconsciously due to the pervasive nature of prejudice in society. This makes automated detection of injustices more challenging which leads to them being often overlooked. We introduce a novel framework that combines knowledge from epistemology to enhance the detection of implicit injustices in text using NLP models to address these complexities and offer explainability. Our empirical study shows how our framework can be applied to effectively detect these injustices. We validate our framework using a human baseline study which mostly agrees with the choice of implicit bias, stereotype, and sentiment. The main feedback from the study was the extended time required to analyze, digest, and decide on each component of our framework. This highlights the importance of our automated framework pipeline that assists users in detecting implicit injustices while offering explainability and reducing time burdens on humans.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
@@
The most basic duty of the media is knowledge sharing. Yet, the tool necessary for wide-spread knowledge sharing is influence. With this influence, the media is able to shape how one will understand the intricacy of a storyline in a news story with little effort. This often results in the use of epistemic biases which involves propositions that are presupposed, entailed, asserted, hedged, or boosted in text (Recasens et al., 2013) to erode or assert a person’s capacity as a knower, leading to framing issues and implicit injustices within the text. Particularly, some of these word choices in text can lead to e.g., epistemic injustices occurring (i.e., when the information is not based on facts or the truth, but more on subjective beliefs). These injustices can affirm or perpetuate stereotypes concerning the subjects involved. Though it has always been a harsh reality with various ringing consequences, in recent years we have publicly witnessed how the affirmation of stereotypes can lead to physical violence, prejudice, and negative self-image (Harrison and Esqueda, 1999; Gover et al., 2020; Kuykendall, 1989). These experiences are harmful and dangerous, explicitly for the victims, but also for all members of our society.
@@
Therefore, we seek to detect instances of testimonial injustice, character injustice, and framing injustice in text. Testimonial injustice occurs when word choices cast modified believability on a statement due to prejudices (e.g., stereotypes) about the sayer (Fricker, 2007). Character assassination is “the deliberate destruction of an individual’s reputation or credibility” (Icks et al., 2019). This often leads to character injustice, a term we define as an unjustified attack on a person’s character that results in an unfair criticism or inaccurate representation of them (e.g., exaggeration or defamation). According to (Entman, 2007) framing bias happens when the use of subjective, one-sided words that reveals the stance of an author occurs. Which means that an individual’s choice from a set of options is influenced more by how the information is worded, rather than by the information itself. We recognise that text, such as those found in news content can be positively or negatively framed to influence the narratives. In this paper, we aim to show how word choices that negatively frame a statement, rather than presenting it with factual or neutral terms, can result in the unfair depiction of certain subjects, thereby contributing to them experiencing these injustices.
@@
With this work, we seek to make room for subjects of text, even in creative writing, to not have their credibility shot or character assumed due to well-known stereotypes which are harmful and unfounded. It can be noted here that toxic or hate speech is often blatant and highly recognisable. Which differs from the objective of our work since here we focus on implicit injustices which are less obvious but could be just as harmful as toxic or hate speech. We pursue this, acknowledging it is achievable to speak about any happening in a factual way that does not seek to cause harm on the subject of the text. Further, we seek to offer explainability towards how and why a writing may cause or promote injustice towards the subject of their text.
@@
Our proposed framework will help with the detection of testimonial, character, and framing injustices. The framework includes a fine-tuned BERT model based on work from (Pryzant et al., 2020b) to automatically tag words associated with epistemic bias from an input text, uses the (Kwon and Gopalan, 2021) CO-STAR model and (Sap et al., 2020) Social Bias Frames (SBF) to generate potential associated stereotypes and the concepts of those stereotypes to the input text, and show when the tagged words (associated with some epistemic bias) or less credibility of a person are correlated with a stereotype which causes injustice. Though we could use examples from various fields (e.g., politics, marketing, medicine, etc…), we will use news media as a use-case throughout this work. Thus, we present the following contributions: (1) We develop a novel framework that uses the results of 3 NLP models to automatically detect character, testimonial, and framing injustices in text, (2) We produce a fine-tuned tagger model to automatically detect epistemic biases, (3) We develop an Interactive User-Interface for journalists and editors to submit text to and receive output and explanations surrounding the potential bias inducing tagged word, (4) We produce empirical evidence showing how epistemic bias can translate to injustices, and (5) We conduct a survey to validate the results of our empirical study and feature selection for fine-tuning the tagger model.
@@
An outcome of this work is to give journalists and editors a tool which will help them easily and quickly notice and avoid testimonial, character, and framing injustices in their work. This will be accomplished by showing users which words they use that produce epistemic bias, the potential stereotype associated with the tagged words and text, offering the user explainability with the help of the stereotype concepts as defined by (Kwon and Gopalan, 2021), and resources to reference literature on the particular epistemic bias type(s) identified in their input text.
2. Background
@@
Many works have established it is difficult for the common person to identify a biased word in a sentence and establish the need for computational agents to take on this charge. Pryzant et al. (2020b) show humans have low ability to detect bias and show humans perform worse than their detection model. Recasens et al. (2013) show that the accuracy of Humans annotators on AMT (amazon mechanical turk) was not more than 37.39% for a single detected biased word. The difficulty arises due to us holding our own biases as facts and lack of education on sentence construction. It has also been observed that when human, expert annotators follow guides that help them detect bias, can more easily find instances of such biases. For example, the editors of the Wiki Neutrality Corpus had to follow the Netural Point of View (NPOV) Guidelines  [^1] https://en.wikipedia.org/wiki/Wikipedia:Neutral_point_of_view.
@@
Great efforts have been put towards identifying potential words in text materials that could encourage epistemic bias (Recasens et al., 2013; Hube and Fetahu, 2019; Pryzant et al., 2020b). Following on from the work of Pryzant et al. (2020b), we fine-tuned their tagger model to automatically tag words associated with epistemic bias from an input text. Identifying words which cause epistemic bias is a step towards awareness of social harms. This begs the question, what do we do with our new found knowledge and awareness? What kinds of implications does our use of these words impact society? What communities are affected by these word choices? These are the questions we explore in this through our empirical study and validation study as discussed in Section 3.
@@
Authors Kwon and Gopalan (2021) have trained a model to detect widely-known stereotypes and the concept of those stereotypes in text materials. We leverage the results of their CO-STAR model and the SBF model (Sap et al., 2020) to offer some explainability of the word choices by the model. Associating a particular text with a stereotype and the concept of that stereotype is a critical step towards awareness of social harms that might cause character or testimonial injustice to a particular individual or group. Lack of identifying the words in a sentence which imply and promote these stereotypes leaves us with the undirected burden and question of: how can we address these harms? Our solution promotes awareness and accountability to directly highlight words that are likely contributing to such harms (using a fine-tuned tagging model), enabling users to recognize and address potentially harmful language more effectively.
@@
Raza et al. (2022) developed a pipeline which takes in news articles, detects and masks words that are biased, and suggests words with more neutral text. Whilst the pipeline and library designed by the authors are very good and useful, they however do not consider the linguistic and epistemic bias features as discussed in Recasens et al. (2013) and used by Pryzant et al. (2020b). Unlike our work discussed in this paper, their pipeline produces no way of distinguishing and highlighting epistemic bias types of any identified potential biased word in the sentence. Also, their framework does not attempt to relate potential stereotypes and stereotype concepts that might be associated with causing injustice to a person or group.
3. Methodology
Hamborg et al. (2019) concluded that various forms of media bias is already analysed in the Social Sciences field and can be implemented in an automated fashion by primarily using Natural Language Processing (NLP) techniques. We leverage research methods using NLP and deep learning whilst also using analysis concepts by researchers from social sciences. Thus, we propose a novel, technical framework (Figure 1) which includes using NLP models for detecting potential epistemic biased words and potential stereotypes along with their concepts, which we semantically link to given sentence(s). From the literature, we found very few models which consider linguistic, epistemic features, or common sense reasoning to detect implicit biases or reasoning about potential stereotypes. This in turn influenced our decision to use three baseline models with promising results (i.e., Tagger Model, Co-Star, and SBF).
Figure 1. Testimonial Injustice Technical Framework
@@ 3. Methodology
Figure 2. Process of Getting Data to the Framework
3.1. Tagger model
@@
The tagger model (see Figure 9 in Appendix A) takes as input sentence(s) (e.g., a headline text, a collection of headlines, etc…), and predicts the probability of each word in the sentence(s) been biased (see Equation 1 and Figure 1). Then it returns the word with the highest probability as the tagged biased word.
(1) $P_{i}=\sigma(b_{i}W^{b}+e_{i}W^{e}+b)$
Where $b_{i}\in R^{b}$ is a word’s $w_{i}^{s}$ semantic meaning and $e_{i}$ are the experts features as proposed by (Recasens et al., 2013) and based on Equation 2.
(2) $e_{i}=R_{e}LU(f_{i}W^{in})$
The detection model itself is a BERT-based neural sequence tagging model that has been fine-tuned to include the expert linguistic and epistemic bias features from Recasens et al. (2013). An example of the expert feature as listed in Table 3 of Recasens et al. (2013) is the part of speech (POS) tag of each word in the sentence. Another example of an expert feature from vetted experts in linguistics is “assertive verbs”. The model leverages the semantic meaning of each word in the given sentence via the BERT (Kenton and Toutanova, 2019) contextualised word vector.
When there are words in a sentence which do not help with understanding, but are excess information these are known to cloud the judgment of the reader while simultaneously giving the reader reason to be more confident in the conclusions they make (Lynch, 2023; Spira, 2011). We hypothesised here that - having too much information (TMI) can be considered from an English linguistic standpoint to be a sentence instance which has more than 2 descriptive words (i.e., adjectives and adverbs) in it that do not add to its understanding but seeks to cloud the judgement of the readers. We manually analysed a few biased and non-biased headlines and saw an excessive use of descriptors (i.e., adjectives, adverbs) in headlines that were biased. We extend the detection model further by including a feature on whether or not the sentence contained TMI. We created methods in python which makes use of the core NLP dependency tree and a tree traversal algorithm to go through sentences starting from the root node, then count the number of adjectives and adverbs in the sentence to determine if there is “TMI” or “no TMI” based on the hypothesis. TMI is not a known indicator of epistemic bias, but we acknowledge it might contribute to causing doubt which leads to injustice. This is why we introduce this new feature to see if it will be a contributing feature for the tagger model to detect epistemic biases. We conducted an ablation study to see the effects of adding this feature, whilst training the tagger model in Section 4.1. Further, we conduct a survey to validate our hypothesis about TMI in Section 4.3.
3.2. CO-STAR Model and Social Bias Frames
@@
Since testimonial injustice occurs because of widely known stereotypes, the outputs of these models will help us inform our users of any potential stereotypes that are exacerbated by their inputted text. The potential stereotype can also be used to determine if a person’s character has been unjustly targeted. This is because stereotypes are not based on the actual truth of the particular person they are attached to and stigmatises the individual. Their presence amidst character assassination is evidence of character injustice. Stereotype concepts will offer explainability as to why a certain word was tagged as being epistemically biased.
We submit news article headlines to the CO-STAR and SBF models and receive an output of 6 potential stereotypes and 3 stereotype concepts the sentence is related to. Semantic similarity describes how closely related two items (e.g., 2 words or 2 sentences) are in terms of meaning. We use semantic similarity to determine how closely related each of the generated stereotype sentences in the list are to the original headline sentence. After obtaining the vector embedding for the sentences, we used a distance metric on the encoded vectors (cosine similarity metric from sentence transformers), to get the distance scores, then we ranked the outputs based on their semantic similarity to the headlines. The closest stereotype and stereotype concept to the sentence is output to the user as the potential stereotype and concept which casts doubt on the subject of the inputted text.
3.3. Lexicon Lookup
Once the tagger model has returned the top tagged word, we proceed to automatically look up and semantically search the tagged word in the epistemic lexicons from the social sciences - to discover the epistemic bias types it is associated with. These lexicons are from the collection of datasets we discuss in the dataset Section 3.5. When a tagged word is not found in the lexicons, we lemmatize said word (considering its context) to find its base word and find the stem word (removing the prefixes and suffixes). We then search for the lemmatized and stemmed words in the lexicons.
3.4. Interactive Interface for Learning about Bias
@@
During the analysis, we first leverage the nltk python library to remove any stop words (i.e., a, the, etc..). We also check the sentence length. We make an assumption here, that if the sentence is less than 3 words, then there is not enough context to analyse the text for potential stereotypes. However, if we have enough context, the content is sent to the epistemic tagger model to find the word with the highest probability of bias (see Figure 3). The conditions from the framework (Figure 1) are then implemented.
In Figure 3 and 4 show screenshots of our UI after a user has submitted a text for analyzing. The word suspected of causing injustice is highlighted with the certainty score, an explanation of why the word was tagged. When the user clicks Show Details they will see more details as displayed in Figure 4.
Figure 3. The interactive interface showing the highlighted tagged word and associated information for a sentence.
Figure 4. UI showing the tagged words, their associated probability for bias, the potential stereotype and stereotype concept and some explainable data created from the lexicon search.
We display the types of associated epistemic bias detected to the user and provide links to resources which explains more on the specific types of biases as shown in Figure 4. See examples of these outputs in table 7 and table 6 of Appendix A.
3.5. Datasets